# D3: Diversity, Difficulty, and Dependability-Aware Data Selection for Sample-Efficient LLM Instruction Tuning

### v1

@@ Abstract
Recent advancements in instruction tuning for large language models (LLMs) suggest that a small, high-quality dataset can significantly equip LLMs with instruction-following capabilities, outperforming large datasets often burdened by quality and redundancy issues. However, the challenge lies in automatically identifying valuable subsets from large datasets to boost both the effectiveness and efficiency of instruction tuning. In this paper, we first establish data selection criteria based on three distinct aspects of data value: diversity, difficulty, and dependability, and then propose the D3 method comprising two key steps of scoring and selection. Specifically, in the scoring step, we define the diversity function to measure sample distinctiveness and introduce the uncertainty-based prediction difficulty to evaluate sample difficulty by mitigating the interference of context-oriented generation diversity. Additionally, we integrate an external LLM for dependability assessment. In the selection step, we formulate the D3 weighted coreset objective, which jointly optimizes three aspects of data value to solve for the most valuable subset. The two steps of D3 can iterate multiple rounds, incorporating feedback to refine the selection focus adaptively. Experiments on three datasets demonstrate the effectiveness of D3 in endowing LLMs with competitive or even superior instruction-following capabilities using less than 10% of the entire dataset.
@@ 1 Introduction
With the success of large language models (LLMs) Brown et al. (2020); Touvron et al. (2023a, b), the application of data-driven AI has been significantly advanced across various industries and scenarios. In the LLM paradigm, large-scale pre-training on massive corpora equips models with knowledge and text generation capabilities based on next-token prediction Naveed et al. (2024). Several post-training techniques Bai et al. (2022); Wang et al. (2024b); Rafailov et al. (2024) could further enhance the model’s performance on downstream tasks. In particular, instruction tuning Ouyang et al. (2022); Taori et al. (2023); Zhang et al. (2024), which fine-tunes models with labeled pairs of instruction and response, plays a crucial role in enabling LLMs’ instruction-following capabilities across various tasks.
@@ 2.1 Instruction Tuning for LLMs
Instruction tuning Brown et al. (2020); Touvron et al. (2023a, b) has emerged as a widely adopted post-training strategy to equip language models with instruction-following capabilities for various downstream tasks. Early research Khashabi et al. (2020); Ye et al. (2021) primarily focused on manually creating task-specific instruction tuning datasets. More recent approaches Honovich et al. (2022); Wang et al. (2022); Xu et al. (2023); Wang et al. (2023); Taori et al. (2023) shift toward constructing instruction tuning datasets by leveraging powerful LLMs to generate data. These studies facilitate the automatic data creation via distilling knowledge from the teacher model, and are also related to the growing field of LLM-based data synthesis Wang et al. (2024a); Long et al. (2024). These studies have laid the foundation for instruction tuning, demonstrating that large task-specific or cross-task instruction tuning datasets can boost the instruction-following performance of language models.
@@ 2.2 Sample-Efficient Tuning
Figure 1: The overall framework of D3 data selection method, including the warm-up step and two key steps of scoring and selection.
@@ 3.3.2 Sample Difficulty
$\mathcal{H}_{t}^{(i)}=-\sum_{i}p(w_{i}\mid x^{(i)},y_{<t}^{(i)})\log(p(w_{i}\mid x^{(i)},y_{<t}^{(i)})),$ (5)
where $\mathcal{H}_{t}^{(i)}$ denotes the entropy of the $t$-th token prediction given $x^{(i)}$ and $y_{<t}^{(i)}$, and $w_{i}$ represent each token in vocabulary.
@@
$d_{2}(z^{(i)})=\text{UDP}(z^{(i)})=\frac{1}{|y^{(i)}|}\sum_{t}\text{UDP}_{t}^{(i)},$ (7)
@@
Algorithm 1 The D3 Data Selection Procedure.
@@
Winning Score (vs. Full) Leaderboard
Dataset Strategy WizardLM (218) Sinstruct (252) Vicuna (80) Koala (180)
@@
Average (1030) AlpacaEval
Alpaca (5%) Full 1.00 1.00 1.00 1.00 1.00 1.00 28.00
PPL 0.20 0.30 0.08 0.16 0.15 0.19 4.34
Random 0.99 0.93 1.09 0.98 0.95 0.97 28.29
@@ 3.3.3 Sample Dependability
The difficulty score based on UPD helps identify samples with greater difficulty that less likely to be caused by CoGD. However, if a sample itself contains factual errors, extraneous symbols, or lacks fluency, it can undermine both the data selection and instruction tuning. To tackle this, we integrate an external teacher model to assess the dependability of samples through prompt-based evaluation. This helps filter out unreliable samples, reducing the risk of incorporating erroneous supervision during instruction tuning. By leveraging the teacher’s instruction-following capability, we obtain a soft dependability score $d_{3}$ within $[0,1]$ for each sample:
$\small d_{3}(z^{(i)})=\frac{\exp(F^{*}(w_{pos}\mid T(z^{(i)})))}{\exp(F^{*}(w_{pos}\mid T(z^{(i)})))+\exp(F^{*}(w_{neg}\mid T(z^{(i)})))},$ (8)
@@ 3.4 D3 Selection Object
$=\operatorname*{arg\,min}_{\mathcal{Q}}\max_{z\in\mathcal{D}}\min_{z^{\prime}\in\mathcal{Q}}d_{2}(z)d_{3}(z)\mathbf{d}_{\psi_{F}}(z,z^{\prime}),$
@@ 4.1 Experimental Setup
  • Random: A subset of data is randomly sampled from the instruction tuning dataset.
@@
To ensure fairness and adherence to the settings and procedure of candidate methods, we use the pre-trained Llama2-7B as the initial model. All methods first undergo two training epochs on a warm-up set, which consists of 1% of the data randomly sampled from the data pool. Subsequently, each strategy selects a data subset of equal size for instruction tuning based on a given selection budget. Samples selected are trained for three epochs to obtain the fine-tuned LLMs. More details on training are provided in the appendix.
@@ 4.3 Main Results
We present the primary comparison results in Table 1. For all three instruction tuning datasets, we allocate a $k=5\%$ selection budget for sample-efficient instruction tuning. To ensure a fair comparison, all strategies select data to reach the budget in one go, except for the Full strategy, which uses all the data. Specifically, we set total round $R=1$ in our D3. Our proposed D3 method demonstrates superior results by effectively selecting a small, yet highly valuable subset of data, enhancing the model’s instruction-following capabilities on both medium and high-quality instruction tuning datasets. It outperforms the Full strategy in both performance and efficiency. In contrast, using perplexity, or loss, as a measure of sample value for data selection guidance fails to account for diversity and difficulty, resulting in considerably poorer performance. The Random strategy, which can be seen as a simple approach that focuses solely on diversity to capture the overall data distribution by random sampling, achieves performance slightly weaker than Full. However, it falls short of outperforming Full due to the lack of more in-depth sample-level evaluation. The IFD method provides a deeper analysis of sample-wise instruction-following difficulty, helping the model enhance performance with less data. However, as it relies solely on a single metric for data selection, it leaves room for further improvement. Besides, the results reported by AlpacaEval show that it is slightly less competitive than the Full strategy on two high-quality instruction tuning datasets.
Table 2: Comparisons of D3 with three ablation strategies. Scores represent the average winning scores across all test test.
@@ 4.4 Performance with Varied Budgets
We systematically examine the impact of selection budgets on various strategies by incrementally increasing the budget. In line with the setup in the main experiment, all strategies select the required data in a single round. Figure LABEL:fig:exp2 presents the winning score on the Alpaca dataset. As the budget increases, the performance of more potent strategies exhibits a trend of initially improving and then declining. Conversely, weaker strategies gradually approach the performance of the Full strategy. This is because selecting more valuable data positively contributes to model instruction tuning. However, once the effective data reaches saturation, incorporating additional data of slightly lower quality or with redundancy leads to diminishing returns, ultimately causing the performance to converge with that of the Full strategy. Specifically, D3 achieves superior performance with the smallest budget, highlighting the effectiveness of our strategy in selecting the most valuable coreset data for sample-efficient instruction tuning. Similarly, the results on the other two high-quality datasets, as reported in the appendix, also empirically validate this. In contrast, other strategies display weaker performance trajectories or slower growth, which further underscores the superiority and effectiveness of our approach.
@@ 4.5 Ablation Study
Table 2 presents a performance of the three ablation strategies. Given that Alpaca is a medium-quality dataset, neglecting diversity causes the model to overly focus on certain challenging and highly dependable samples, resulting in a significant performance drop. In the case of the other two high-quality datasets, most samples have dependability scores close to 1, making D3-w/o2, to some extent, a strategy that only considers diversity. As a result, its performance is similar to but slightly better than the Random strategy shown in the main results. The performance impact of removing dependability is more pronounced for Alpaca due to the quality difference of datasets. These validate the necessity of comprehensively considering all three criteria in data selection.

### v2

@@ Abstract
Recent advancements in instruction tuning for large language models (LLMs) suggest that a small, high-quality dataset can significantly equip LLMs with instruction-following capabilities, outperforming large datasets often burdened by quality and redundancy issues. However, the challenge lies in automatically identifying valuable subsets from large datasets to boost both the effectiveness and efficiency of instruction tuning. In this paper, we first establish data selection criteria based on three distinct aspects of data value: diversity, difficulty, and dependability, and then propose the D3 method comprising two key steps of scoring and selection. Specifically, in the scoring step, we define the diversity function to measure sample distinctiveness and introduce the uncertainty-based prediction difficulty to evaluate sample difficulty by mitigating the interference of context-oriented generation diversity. Additionally, we integrate an external LLM for dependability assessment. In the selection step, we formulate the D3 weighted coreset objective, which jointly optimizes three aspects of data value to solve for the most valuable subset. The two steps of D3 can iterate multiple rounds, incorporating feedback to refine the selection focus adaptively. Experiments on both public datasets and the real-world Taobao Live application demonstrate the effectiveness of D3 in endowing LLMs with competitive or even superior instruction-following capabilities using less than 10% of the entire dataset.
@@ 1 Introduction
With the success of large language models (LLMs) Brown et al. (2020); Touvron et al. (2023), the application of data-driven AI has been significantly advanced across various industries and scenarios. In the LLM paradigm, large-scale pre-training on massive corpora equips models with knowledge and text generation capabilities based on next-token prediction Naveed et al. (2024). Moreover, several post-training methods and techniques Bai et al. (2022); Wang et al. (2024b); Rafailov et al. (2024) could further enhance the model’s alignment and performance on downstream tasks. In particular, instruction tuning Ouyang et al. (2022); Taori et al. (2023), which fine-tunes models with labeled pairs of instruction and response, plays a crucial role in enabling LLMs’ instruction-following capabilities across various tasks.
@@ 2.1 Instruction Tuning for LLMs
Instruction tuning Brown et al. (2020); Touvron et al. (2023) has emerged as a widely adopted post-training strategy to equip language models with instruction-following capabilities for various downstream tasks. Early research Khashabi et al. (2020); Ye et al. (2021) primarily focused on manually creating task-specific instruction tuning datasets. More recent approaches Honovich et al. (2022); Wang et al. (2022); Xu et al. (2023); Wang et al. (2023); Taori et al. (2023) shift toward constructing instruction tuning datasets by leveraging powerful LLMs to generate data. These studies facilitate the automatic data creation via distilling knowledge from the teacher model, and are also related to the growing field of LLM-based data synthesis Wang et al. (2024a); Long et al. (2024). These studies have laid the foundation for instruction tuning, demonstrating that large task-specific or cross-task instruction tuning datasets can boost the instruction-following performance of language models.
@@ 2.2 Sample-Efficient Tuning
Figure 1: The overall framework of D3 data selection method, including the warm-up and two key steps of scoring and selection.
@@ 3.3.2 Sample Difficulty
$\mathcal{H}_{t}^{(i)}=-\sum_{j}p(w_{j}\mid x^{(i)},y_{<t}^{(i)})\log(p(w_{j}\mid x^{(i)},y_{<t}^{(i)})),$ (5)
where $\mathcal{H}_{t}^{(i)}$ is the entropy of the $t$-th token prediction given $x^{(i)}$ and $y_{<t}^{(i)}$, and $w_{j}$ denotes each token in the vocabulary.
@@
$d_{2}(z^{(i)})=\text{UPD}(z^{(i)})=\frac{1}{|y^{(i)}|}\sum_{t}\text{UPD}_{t}^{(i)},$ (7)
@@
Algorithm 1 The D3 Data Selection and Fine-tuning.
@@
Winning Score (budget $k$ vs. Full) Leaderboard
Dataset
(budget $k$)
Selection
Strategy
WizardLM
Sinstruct
Vicuna
Koala
@@
Average
AlpacaEval
(budget $k$ vs. Full)
Alpaca (5%) PPL 0.20 0.30 0.08 0.16 0.15 0.19   4.34 28.00
RAND 0.99 0.93 1.09 0.98 0.95 0.97 28.29
@@ 3.3.3 Sample Dependability
The difficulty score based on UPD helps identify samples with greater difficulty that less likely to be caused by CoGD. However, if a sample itself contains factual errors, extraneous symbols, or lacks fluency, it can undermine both the data selection and instruction tuning. To tackle this, we integrate an external teacher model to assess the dependability of samples through prompt-based evaluation. This helps filter out unreliable samples, reducing the risk of incorporating erroneous supervision during instruction tuning. By leveraging the teacher’s instruction-following capability, we obtain a soft dependability score $d_{3}$ within $(0,1)$ for each sample:
$d_{3}(z^{(i)})=\frac{\exp(F^{*}(w_{pos}\mid T(z^{(i)})))}{\exp(F^{*}(w_{pos}\mid T(z^{(i)})))+\exp(F^{*}(w_{neg}\mid T(z^{(i)})))},$ (8)
@@ 3.4 D3 Selection Objective
$=\operatorname*{arg\,min}_{\mathcal{Q}}\max_{z\in\mathcal{D}}\min_{z^{\prime}\in\mathcal{Q}}d_{2}(z)d_{3}(z)\mathbf{d}_{\psi_{F}}(z,z^{\prime}),$ (9)
@@ 4 Empirical Study
Table 2: Pairwise comparisons on real-world Taobao Live dataset across five evaluation dimensions of interest: Satisfaction, Presentation Skill, Conversion, Spoken Fluency, and Overall Performance.
Winning Score (budget 10% vs. Full)
Strategy Satis. Pres. Conv. Spoke. Overall
PPL 0.87 0.84 0.81 0.89 0.82
RAND 0.90 0.91 0.92 0.89 0.93
IFD 0.77 0.74 0.76 0.82 0.71
D3 1.07 1.07 1.04 1.06 1.11
@@ 4.1 Experimental Setup
Real-world Dataset. We further evaluate our method in a real-world setting within Taobao Live, where the language model is fine-tuned using collected live-streaming data to generate high-quality sales-pitch scripts tailored to product information, thereby empowering streamers to improve their performance. We utilize 64,881 collected samples for training, with a random selection of 300 reserved for testing. A pairwise evaluation of generated scripts is conducted across five evaluation dimensions of interest in the task context.
@@
  • RAND: A subset of training data is randomly sampled from the original instruction tuning dataset.
@@
To ensure fairness and adherence to the settings and procedure of candidate methods, we use the pre-trained Llama2-7B as the initial model. All methods first undergo two training epochs on a warm-up set, which consists of 1% of the data randomly sampled from the data pool. Subsequently, each strategy selects a data subset of equal size for instruction tuning based on a given selection budget. Samples selected are trained for three epochs to obtain the fine-tuned LLMs.
@@ 4.3 Main Results
We present the primary comparison results in Table 1. For all three instruction tuning datasets, we allocate a $k=5\%$ selection budget for sample-efficient instruction tuning. To ensure a fair comparison, all strategies select data to reach the budget in one go, except for the Full strategy, which uses all the data. Specifically, we set total round $R=1$ in our D3. Our proposed D3 method demonstrates superior results by effectively selecting a small, yet highly valuable subset of data, enhancing the model’s instruction-following capabilities on both medium and high-quality instruction tuning datasets. It outperforms the Full strategy in both performance and efficiency. In contrast, using perplexity, or loss, as a measure of sample value for data selection guidance fails to account for diversity and difficulty, resulting in considerably poorer performance. The RAND strategy, which can be seen as a simple approach that focuses solely on diversity to capture the overall data distribution by random sampling, achieves performance slightly weaker than Full. However, it falls short of outperforming Full due to the lack of more in-depth sample-level evaluation. The IFD method provides a deeper analysis of sample-wise instruction-following difficulty, helping the model enhance performance with less data. However, as it relies solely on a single metric for data selection, it leaves room for further improvement. Besides, the results reported by AlpacaEval show that it is slightly less competitive than the Full strategy on two high-quality instruction tuning datasets. The performance comparisons on the real-world dataset are presented in Table 2, under a fixed selection budget $k=10\%$, further demonstrating the superiority of our method.
Table 3: Comparisons of D3 with three ablation strategies. Scores represent the average winning scores across all test sets.
@@ 4.4 Performance with Varied Budgets
We systematically examine the impact of selection budgets on various strategies by incrementally increasing the budget. In line with the setup in the main experiment, all strategies select the required data in a single round. Figure 3 presents the winning score on the Alpaca dataset. As the budget increases, the performance of more potent strategies exhibits a trend of initially improving and then declining. Conversely, weaker strategies gradually approach the performance of the Full strategy. This is because selecting more valuable data positively contributes to model instruction tuning. However, once the effective data reaches saturation, incorporating additional data of slightly lower quality or with redundancy leads to diminishing returns, ultimately causing the performance to converge with that of the Full strategy. Specifically, D3 achieves superior performance with the smallest budget, highlighting the effectiveness of our strategy in selecting the most valuable coreset data for sample-efficient instruction tuning. Similarly, the results on the other two high-quality datasets, as reported in the appendix, also empirically validate this. In contrast, other strategies display weaker performance trajectories or slower growth, which further underscores the superiority and effectiveness of our approach.
@@ 4.5 Ablation Study
Table 3 shows the performance of the three ablation strategies. Given that Alpaca is a medium-quality dataset, neglecting diversity causes the model to overly focus on certain challenging and highly dependable samples, resulting in a significant performance drop. In the case of the other two high-quality datasets, most samples have dependability scores close to 1, making D3-w/o2, to some extent, a strategy that only considers diversity. As a result, its performance is similar to but slightly better than the RAND strategy shown in the main results. The performance impact of removing dependability is more pronounced for Alpaca due to the quality difference of datasets. These validate the necessity of comprehensively considering all three criteria in data selection.
@@ 5 Conclusion
Appendix A Appendix
This appendix provides supplementary details to support the main text, including further descriptions of the methodology and additional implementation details and results in experiments.
A.1 Methodological Supplement
A.1.1 Sample Diversity
As described in the main text, the sample diversity function we defined depends on the calculation between the embeddings of different samples. The sample embeddings are obtained from the LLM to be fine-tuned, rather than introducing external models, thereby providing an unbiased representation of the relationships between samples within the current LLM’s feature space. Due to the decoder-based transformer architecture of current mainstream LLMs, it is hard to directly obtain a semantic embedding for each sample. Therefore, as shown in Algorithm 2, we average the final-layer hidden features of each token in the sample to obtain a vector as its embedding.
Algorithm 2 The Sample Embedding Calculation.
1:  Input: Sample $z=\{x,y\}$, LLM $F$.
2:  Procedure:
3:  $\hat{x},\hat{y}\leftarrow\mathrm{tokenization(x,y)}$
4:  for $t:=0\rightarrow|\hat{y}|$ do
5:     $\mathrm{hidden}_{t}\leftarrow F_{\mathrm{last\_hidden}}(\hat{x},\hat{y}_{<t})$
6:  end for
7:  $\mathrm{embedding}\leftarrow\frac{1}{|\hat{y}|}\sum_{t=0}^{|\hat{y}|}\mathrm{hidden}_{t}$
8:  return  $\mathrm{embedding}$
A.1.2 Sample Dependability
We introduce an external teacher LLM to evaluate sample-model dependability. We generate a prompt for each sample and input it to the teacher model to obtain an assessment of the sample’s dependability. Specifically, the prompt template, as shown in Figure 5, presents dependability assessment criteria in the form of instructions. These criteria can be tailored to different tasks by customizing specific evaluation preferences. The teacher model then outputs an evaluation score of 0 (bad) or 1 (good) on a new line, corresponding to negative and positive tokens $w_{neg}$ and $w_{pos}$, respectively. For evaluation efficiency, we limit the maximum number of tokens generated by the teacher model to 1 and truncate subsequent explanatory content related to the score. We compute the soft scores as sample dependability using the logits output by the teacher model on these two tokens.
Figure 5: The prompt template used to evaluation the sample depenability.
Sample Dependability Evaluation Prompt
You are tasked with evaluating the sample based on whether it meets all the following criteria:
- Fluency: Is the [Response] coherent and free from irrelevant content or nonsensical marks?
- Accuracy: Does the [Response] correctly answer the [Query] without providing false information?
- Clarity: Is the [Response] clear and logically structured?
[Query]
{instruction}
[Response]
{response}
Please review the sample and assign a quality score of 0 (bad) or 1 (good).
Quality score (0 or 1):
$d_{3}(z)=\frac{\exp(F^{*}(w_{pos}\mid T(z)))}{\exp(F^{*}(w_{pos}\mid T(z)))+\exp(F^{*}(w_{neg}\mid T(z)))},$ (11)
where $F^{*}$ represents the teacher model, $T(z)$ is the evaluation prompt generated based on sample $z$, and $F^{*}(w|T)$ refers to the logit of the next token $w$ given the inputs $T(z)$.
A.2 Experimental Supplement
A.2.1 Evaluation Details
In experiments, we employ two evaluation methods: pairwise evaluation and the AlpacaEval Leaderboard Li et al. (2023). The pairwise evaluation employs a more powerful model as a judge to assess the quality of two different responses to the same instruction. Specifically, we construct the pairwise evaluation prompt as shown in Figure 6 and pass it to the judge LLM for evaluation. The judge model used for evaluation inference in our experiments is the locally deployable Qwen2.5-32B-Instruct Team (2024); Yang et al. (2024), which possesses extensive knowledge and strong instruction-following capabilities. The test datasets involved in pairwise evalutation are WizardLM Xu et al. (2023), Self-instruct Wang et al. (2023), VicunaChiang et al. (2023), Koala Vu et al. (2023), and LIMAZhou et al. (2023). They consist of 218, 252, 80, 180, 300 human-curated instructions, respectively, spanning domains such as mathematics, coding, writing, knowledge, and computer science. For the AlpacaEval Leaderboard, we deployed the official AlpacaEval repository [^1] https://github.com/tatsu-lab/alpaca_eval locally and used the GPT-4o API for evaluation. Additionally, the widely used Open LLM Leaderboard benchmark could not be deployed locally for evaluation in this experiment due to the absence of certain original test datasets in the repository. Given the large scale and high credibility of the two evaluation methods described above, this leaderboard was temporarily excluded but may be incorporated in future evaluations.
Figure 6: The prompt template used for pairwise evaluation of the model response quality.
Pairwise Evaluation Prompt for Response
System Prompt
You are a helpful and precise assistant for checking the quality of the answer.
User Prompt
[Question]
{instruction}
[The Start of Assistant 1’s Answer]
{response 1}
[The End of Assistant 1’s Answer]
[The Start of Assistant 2’s Answer]
{response 2}
[The End of Assistant 2’s Answer]
We would like to request your feedback on the performance of two AI assistants in response to the user question displayed above.
Please rate the helpfulness, relevance, accuracy, level of details of their responses. Each assistant receives an overall score on a scale of 1 to 10, where a higher score indicates better overall performance. Please first output a single line containing only two values indicating the scores for Assistant 1 and 2, respectively. The two scores are separated by a space. In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias and ensuring that the order in which the responses were presented does not affect your judgment.
A.2.2 Implementation Details
In experiments, we use the pre-trained Llama2-7B Touvron et al. (2023) as the initial model. All methods begin with two training epochs on a warm-up set, comprising 1% of the data randomly sampled from the data pool. Next, each strategy selects a data subset of equal size to reach the given selection budget for instruction tuning. The selected samples are then trained for three epochs from the warm-up model to obtain the final fine-tuned LLMs. The model training and inference are implemented based on the Llama-Factory Zheng et al. (2024) framework. We use the AdamW optimizer with a batch size of 128. The maximum input length for model training is 2048, while the maximum output length for inference is 512. The initial learning rate for the warm-up step is set to 1e-5, with a starting rate of 2e-5 for training on selected data. A cosine annealing learning rate schedule is applied during the process. In the multi-round selection of D3(MR), each round involves training for 2 epochs on the selected data in last round. Finally, the warm-up model is trained using the same setting as described above, based on all the selected data from each round, to obtain the final model. The sample dependability is calculated using the Qwen2.5-14B-Instruct Yang et al. (2024); Team (2024) model. For the real-world dataset, we use Qwen2.5-7B as the initial model, with other settings consistent with the above discussion.
A.2.3 Additional Experimental Results
In this section, we provide the detailed results for each experiment.
A.2.4 Performance with Varied Budgets
The results with varied budgets of different strategies are presented in Table 4, 5 and  6 for the Alpaca, AlpacaGPT4 and WizardLM datasets, respectively.
A.2.5 Ablation Study
The detailed comparison results of the ablation strategies are provided in Table 7.
A.2.6 Selection with Multi-Round
The detailed comparison results of D3 and D3(MR) are presented in Table 8.
Table 4: Performance with varied selection budgets on the Alpaca dataset.
Winning Score (vs Full) Leaderboard
Strategy Budget WizardLM Sinstruct Vicuna Koala LIMA Average AlpacaEval
Full 1.000 1.00 1.00 1.00 1.00 1.00 1.00 28.00
D3 0.025 1.22 1.14 1.33 1.26 1.38 1.26 40.49
0.050 1.22 1.16 1.21 1.33 1.37 1.27 42.01
0.075 1.31 1.20 1.23 1.21 1.34 1.27 43.62
0.100 1.29 1.23 1.37 1.33 1.36 1.31 43.12
0.125 1.25 1.17 1.18 1.28 1.31 1.25 42.06
0.150 1.21 1.21 1.17 1.24 1.27 1.23 40.04
PPL 0.025 0.11 0.19 0.04 0.13 0.07 0.12 -
0.050 0.20 0.30 0.08 0.16 0.15 0.19 4.34
0.075 0.29 0.38 0.09 0.28 0.17 0.26 -
0.100 0.42 0.46 0.19 0.43 0.33 0.39 9.27
0.125 0.63 0.61 0.65 0.61 0.48 0.58 -
0.150 0.71 0.6 0.64 0.66 0.68 0.66 -
Random 0.025 0.84 0.79 0.91 0.90 0.93 0.87 26.69
0.050 0.99 0.93 1.09 0.98 0.95 0.97 28.29
0.075 0.95 0.90 0.99 0.95 0.99 0.95 28.09
0.100 0.88 0.96 0.91 0.94 0.93 0.93 28.24
0.125 0.98 0.89 1.03 0.98 1.03 0.98 29.87
0.150 0.98 0.98 0.94 1.01 1.04 1.00 30.35
IFD 0.025 1.13 0.89 1.48 1.1 1.28 1.14 -
0.050 1.15 1.01 1.38 1.13 1.28 1.17 36.50
0.075 1.08 1.14 1.49 1.24 1.30 1.22 -
0.100 1.14 1.11 1.46 1.27 1.27 1.22 39.49
0.125 1.22 1.13 1.46 1.25 1.25 1.23 -
0.150 1.11 1.12 1.35 1.16 1.23 1.17 -
Table 5: Performance with varied selection budgets on the AlpacaGPT4 dataset.
Winning Score (vs Full) Leaderboard
Strategy Budget WizardLM Sinstruct Vicuna Koala LIMA Average AlpacaEval
Full 1.00 1.00 1.00 1.00 1.00 1.00 1.00 65.68
D3 0.05 1.13 1.19 1.07 1.24 1.22 1.19 71.26
0.10 1.11 1.20 1.06 1.28 1.22 1.19 72.02
0.15 1.09 1.23 1.06 1.18 1.19 1.17 68.28
0.20 1.09 1.12 1.06 1.24 1.19 1.15 71.27
PPL 0.05 0.29 0.33 0.11 0.36 0.20 0.27 20.92
0.10 0.62 0.69 0.51 0.65 0.66 0.65 -
0.15 0.70 0.66 0.68 0.62 0.74 0.69 -
0.20 0.77 0.79 0.71 0.87 0.87 0.82 -
Random 0.05 0.85 0.86 0.83 0.92 0.98 0.90 60.93
0.10 0.89 1.05 0.99 1.00 0.86 0.95 -
0.15 0.89 1.02 0.91 1.04 0.95 0.97 -
0.20 0.88 0.93 0.89 1.08 0.94 0.95 -
IFD 0.05 1.10 1.08 1.00 1.11 1.15 1.10 64.53
0.10 1.08 1.17 1.06 1.15 1.25 1.16 -
0.15 1.08 1.13 1.17 1.19 1.29 1.18 -
0.20 1.11 1.16 1.18 1.21 1.22 1.18 -
Table 6: Performance with varied selection budgets on the WizardLM dataset.
Winning Score (vs Full) Leaderboard
Strategy Budget WizardLM Sinstruct Vicuna Koala LIMA Average AlpacaEval
Full 1.00 1.00 1.00 1.00 1.00 1.00 1.00 58.42
D3 0.05 0.99 1.14 1.21 1.07 1.17 1.11 63.09
0.10 1.02 1.16 1.15 1.11 1.12 1.11 63.28
0.15 0.97 1.12 1.24 1.10 1.16 1.11 61.00
0.20 0.96 1.12 1.24 1.15 1.12 1.10 62.21
PPL 0.05 0.40 0.52 0.60 0.45 0.48 0.48 28.49
0.10 0.63 0.71 0.94 0.66 0.74 0.71 -
0.15 0.62 0.75 0.93 0.79 0.73 0.74 -
0.20 0.74 0.80 1.06 0.84 0.82 0.82 -
Random 0.05 0.92 0.81 1.08 0.79 0.94 0.89 54.48
0.10 0.93 0.87 1.13 0.98 0.91 0.93 -
0.15 0.91 0.90 1.01 0.93 0.93 0.92 -
0.20 0.92 0.95 1.14 0.98 0.95 0.96 -
IFD 0.05 0.93 0.93 1.28 0.99 1.11 1.02 54.94
0.10 0.97 1.02 1.17 1.03 1.06 1.03 -
0.15 1.01 0.98 1.18 1.04 1.10 1.05 -
0.20 0.97 1.05 1.24 1.10 1.06 1.06 -
Table 7: Detailed results of ablation strategies.
Winning Score (vs. Full)
Dataset Strategy WizardLM (218) Sinstruct (252) Vicuna (80) Koala (180) LIMA (300) Average
Alpaca (5%) D3-w/o1 0.40 0.52 0.21 0.35 0.30 0.38
D3-w/o2 1.10 1.09 1.02 1.07 1.09 1.08
D3-w/o3 1.10 1.04 1.10 1.14 1.19 1.12
D3 1.22 1.16 1.21 1.33 1.37 1.27
AlpacaGPT4 (5%) D3-w/o1 0.84 0.82 0.79 0.84 0.82 0.83
D3-w/o2 0.99 1.07 0.99 1.02 1.03 1.03
D3-w/o3 1.11 1.13 1.14 1.10 1.13 1.12
D3 1.13 1.19 1.07 1.24 1.22 1.17
WizardLM (5%) D3-w/o1 0.76 0.85 0.92 0.81 0.93 0.85
D3-w/o2 0.99 1.02 1.03 1.02 1.01 1.01
D3-w/o3 1.05 1.01 1.20 1.01 1.09 1.06
D3 0.99 1.14 1.21 1.07 1.17 1.11
Table 8: Detailed comparision result of D3 and D3(MR) for three instruction tuning datasets.
Winning Score (vs. Full) Leaderboard
WizardLM Sinstruct Vicuna Koala LIMA Average AlpacaEval
Dataset Budget D3 MR D3 MR D3 MR D3 MR D3 MR D3 MR D3 MR
Alpaca 0.025 1.22 1.22 1.14 1.14 1.33 1.33 1.26 1.26 1.38 1.38 40.49 40.49 40.49 40.49
0.050 1.22 1.27 1.16 1.16 1.21 1.38 1.33 1.34 1.37 1.43 42.01 42.88 42.01 42.88
0.075 1.31 1.33 1.20 1.25 1.23 1.34 1.21 1.34 1.34 1.43 43.62 44.12 43.62 44.12
0.100 1.29 1.42 1.23 1.25 1.37 1.44 1.33 1.37 1.36 1.38 43.12 45.09 43.12 45.09
0.125 1.25 1.31 1.17 1.23 1.18 1.39 1.28 1.39 1.31 1.36 42.06 40.71 42.06 40.71
0.150 1.21 1.29 1.21 1.25 1.17 1.27 1.24 1.31 1.27 1.34 40.04 44.08 40.04 44.08
Alpaca GPT4 0.05 1.13 1.13 1.19 1.19 1.07 1.07 1.24 1.24 1.22 1.22 1.19 1.19 71.26 71.26
0.10 1.11 1.09 1.20 1.21 1.06 1.16 1.28 1.27 1.22 1.21 1.19 1.19 72.02 68.84
0.15 1.09 1.06 1.23 1.13 1.06 1.13 1.18 1.23 1.19 1.18 1.17 1.15 68.28 70.80
0.20 1.09 1.08 1.12 1.17 1.06 1.07 1.24 1.11 1.19 1.17 1.15 1.13 71.27 70.28
WizardLM 0.05 0.99 0.99 1.14 1.14 1.21 1.21 1.07 1.07 1.17 1.17 1.11 1.11 63.09 63.09
0.10 1.02 1.06 1.16 1.16 1.15 1.18 1.11 1.06 1.12 1.19 1.11 1.13 63.28 63.52
0.15 0.97 0.97 1.12 1.16 1.24 1.18 1.10 1.13 1.16 1.19 1.11 1.12 61.00 63.50
0.20 0.96 0.98 1.12 1.14 1.24 1.23 1.15 1.12 1.12 1.17 1.10 1.12 62.21 64.30
Acknowledgments
This research was supported by the Key Program of Jiangsu Science Foundation (BK20243012), the Jiangsu Science Foundation (BG2024036), and the Fundamental Research Funds for the Central Universities (022114380023).
Contribution Statement
Jia Zhang and Yao Liu contribute equally as co-first authors. Yi Liu and Lan-Zhe Guo serve as co-corresponding authors.